40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 19 23:34:52 2012, kiwamu, Update, IOO, PMC realignment and HEPA 
    Reply  Mon Jan 23 15:43:47 2012, Jenne, Update, IOO, PMC realignment 
       Reply  Mon Jan 23 23:12:00 2012, kiwamu, Update, IOO, PMC realignment 
          Reply  Mon Jan 30 08:17:06 2012, steve, Update, IOO, PMC  sickPMC.png
             Reply  Tue Jan 31 08:44:10 2012, steve, Update, IOO, laser shuts down laseroff.png
                Reply  Tue Jan 31 14:58:30 2012, kiwamu, Update, IOO, laser shuts down 
                   Reply  Tue Jan 31 17:13:49 2012, steve, Update, IOO, only the PSL laser is off 
                      Reply  Wed Feb 1 17:00:57 2012, steve, Update, IOO, laser is back ON P1080525.JPGP1080514.JPGP1080518.JPG
                         Reply  Fri Feb 3 14:47:51 2012, steve, Update, PEM, laser interlock drawing 
Message ID: 6236     Entry time: Mon Jan 30 08:17:06 2012     In reply to: 6218     Reply to this: 6239
Author: steve 
Type: Update 
Category: IOO 
Subject: PMC  

Quote:

I have realigned the beam pointing to PMC. The transmitted light increased from 0.74 to 0.83.

The misalignment was mainly in pitch.

 The PMC pointing has changed, so MC is resonating in high order modes.

Attachment 1: sickPMC.png  55 kB  | Show | Show all
ELOG V3.1.3-